OpenStudy (anonymous):

Which of the following is the formula for this geometric sequence? 1, -3, 9, -27,...

OpenStudy (anonymous):

an = -1(3)n-1 an = 1(-3)n-1 an = 3(-1)n-1

OpenStudy (anonymous):

Think: What is a geometric sequence. It's a sequence of the form ar^n. divide any term by the previous term to find r, then use the first term as a.

OpenStudy (anonymous):

That will get you the answer.

OpenStudy (jhannybean):

Geometric sequence : \[\large a_n = a_1r^{n-1}\]\[\large r = -3 ... \frac{-3}{1}=-3 \ , \ \frac{9}{-3} = -3 \ , \ \frac{-27}{9} = -3\]\[\large a_1 = 1\]\[\large a_n = (1)(-3)^{n-1}\]